OpenStudy (anonymous):

log of 6 (x+4) + log of 6 (x-1) = 2

OpenStudy (nikvist):

\[\log_6(x+4)+\log_6(x-1)=2\quad,\quad x>1\] \[\log_6(x+4)(x-1)=2\quad\Rightarrow\quad (x+4)(x-1)=6^2=36\] \[\Rightarrow\quad x^2+3x-40=(x+8)(x-5)=0\quad\Rightarrow\quad x=5\]